If the perimeter of a rectangular house is 44 yards and the length is 36 feet. What is the width of the house?
Answer by AnlytcPhil(1116)

>>...If the perimeter of a rectangular house is 44 yards
and the length is 36 feet. What is the width of the house?
We can either change yards to feet, or feet to yards...<<
1. Changing yards to feet.
Perimeter = 44 yards = 44(3 feet) = 132 feet
Perimeter = 2(Length + Width)
P = 2(L + W)
132 = 2(36 + W)
132 = 72 + 2W
132 - 72 = 2W
60 = 2W
30 = W
So the width is 30 feet.
------------------------------------------
2. Changing feet to yards.
Length = 36 feet = 36(1/3 yard) = 12 yards
Perimeter = 2(Length + Width)
P = 2(L + W)
44 = 2(12 + W)
44 = 24 + 2W
44 - 24 = 2W
20 = 2W
10 = W
So the width is 10 yards.
----------------------------------------
10 yards and 30 feet are the same length.
Edwin